OpenStudy (anonymous):

If you buy a school raffle ticket for $1, your overall probability of winning at least one of the prizes is 0.03. You can improve your odds of winning something by buying two tickets. What is the probability that one or the other of your tickets will win something?(Round your answer to the nearest ten-thousandth.)

OpenStudy (anonymous):

In first case: Chance = 3/100 = 1/(100/3) Second case: Chance = 2/(100/3+1) = 0.58